In the lab, Charmaine has two solutions that contain alcohol and is mixing them with each other. She uses 200 milliliters less of Solution A than Solution B. Solution A is 19% alcohol and Solution B is 11% alcohol. How many milliliters of Solution A does she use, if the resulting mixture has 292 milliliters of pure alcohol?
Answer by stanbon(75887)

She uses 200 milliliters less of Solution A than Solution B. Solution A is 19% alcohol and Solution B is 11% alcohol. How many milliliters of Solution A does she use, if the resulting mixture has 292 milliliters of pure alcohol?
----
Equations:
A = B-200 ml
---
alcohol + alcohol = alcohol
0.19A + 0.11B = 292
----
Substitute for "A" and solve for "B":
0.19(B-200) + 0.11B = 292
---
Multiply thru by 100:
19B - 3800 +11B = 29200
30B = 33000
B = 1100 ml
----
Solve for "A":
A = B - 200
A = 900 ml
=======================
Cheers,
Stan H.
=======================